Stephanie Corporation sells a single product. Budgeted sales for the year are anticipated to be 700,000 units, estimated beginning inventory is 108,000 units, and desired ending inventory is 81,000 units. The quantities of direct materials expected to be used for each unit of finished product are given below.

Material A: 0.50 lb. per unit @ $0.60 per pound
Material B: 1.00 lb. per unit @ $1.76 per pound
Material C: 1.20 lb. per unit @ $0.87 per pound

The dollar amount of Material A used in production during the year is

a.$702,612

b.$1,184,480

c.$201,900

d.$210,000
Answer #1

Answer: $ 201,900

Units Produced = Ending invneotry + sales - Beginning Invneotry

Units Prodcued =81,000+700,000-108,000= 673,000

Prodcution Cost of MAterial A = 673000*0.50*0.60 = $ 201,900 (Answer)